1/1

Михаил Юрьевич Лермонтов,

Sail

tradukita de A. Z. Foreman

A sail drifts white and on its own Amid the light blue ocean haze. What does it seek in distant country? What made it leave its native bays?

Great billows play. High winds are whistling Down at the bending, creaking mast Oh! This one seeks no happy ending And does not flee a happy past.

Beneath, a brighter stream than azure. Above, the golden sunray flows Yet this one, restive...quests for tempests As if in tempests were repose.

Traduko de la Rusa poemo “Парус” de Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) en la Anglan de A. Z. Foreman.

Arg-545-2269 (2015-01-28 19:03:20)

I, Manfred Retzlaff, found this poem-translation in http: // poemsintranslation. blogspot. e/ 2009/ 10/ lermontov-sail-from-russian. html .

http://www.poezio.net/version?poem-id=545&version-id=,2269